OpenStudy (anonymous):

Will Give Medal For Correct Answer Please Show Work Solve: x+12/x+4 = x/x+8

OpenStudy (anonymous):

@skullpatrol can u help?

OpenStudy (skullpatrol):

Can you think of any restrictions on this equation?

OpenStudy (anonymous):

restrictions?

OpenStudy (skullpatrol):

Yes, restrictions are limits on values that the variable can have.

OpenStudy (agent0smith):

\[\Large \frac{ x+12 }{ x+4 } = \frac{ x }{ x+8 }\] yes...?

OpenStudy (anonymous):

@skullpatrol yea, -4, 8, 6, -6 @agent0smith yup thats it

OpenStudy (skullpatrol):

Why did you say -4?

OpenStudy (anonymous):

the series of numbers are the restrictions

OpenStudy (skullpatrol):

Can you give a reason why you chose those particular numbers?

OpenStudy (anonymous):

they were pre chosen for me

OpenStudy (agent0smith):

@skullpatrol those are likely the answer choices. @IanK16. he's asking what values of x will make the equation have a denominator of zero.

OpenStudy (anonymous):

right and i need help with that

OpenStudy (agent0smith):

\[\Large \frac{ x+12 }{ x+4 } \] let's look at the left side first. What value of x will make the denominator (the bottom part) zero? ie... x+4 = 0

OpenStudy (agent0smith):

because... can you divide by zero?

OpenStudy (anonymous):

-4 will make it zero

OpenStudy (agent0smith):

Correct. Now look at the right side of \[\Large \frac{ x+12 }{ x+4 } = \frac{ x }{ x+8 }\] what value of x will give a denominator of zero on the right?

OpenStudy (anonymous):

a -8 will

OpenStudy (agent0smith):

Correct. So those are our "excluded" values, x=-4 and x=-8, in other words, x cannot be equal to those (they aren't solutions). Now you have to solve it, keeping in mind if you get x=-4 or x=-8, they aren't valid solutions.

OpenStudy (anonymous):

ok so what exactly would be the first step?

OpenStudy (agent0smith):

Cross multiply first... do you know how to do that? \[\Large \frac{( x+12) }{( x+4) } = \frac{ x }{ (x+8) } \]

OpenStudy (anonymous):

yes i do, im working on it right now

OpenStudy (anonymous):

\[\frac{ 21x ^{2}+96 }{x ^{2}+4x}\]

OpenStudy (skullpatrol):

@IanK16. What happened to the equals sign?

OpenStudy (agent0smith):

you should get \[\large (x+12)(x+8) = x( x+4)\]

OpenStudy (anonymous):

when i cross multiply i should get that?

OpenStudy (agent0smith):

Yes, cross multiply like this

OpenStudy (anonymous):

right and the answer above is what i got

OpenStudy (agent0smith):

Show us how you got \[\frac{ 21x ^{2}+96 }{x ^{2}+4x}\]

OpenStudy (agent0smith):

Note that in my picture, the equals sign stays where it is... http://assets.openstudy.com/updates/attachments/5170675ee4b0012aba7e2fea-agent0smith-1366325392859-crossmultiply.png

OpenStudy (anonymous):

i cross multiplied. (x+12)(x+8) then (x+4)(x)

OpenStudy (skullpatrol):

Like I said @IanK16. What happened to the equals sign?

OpenStudy (skullpatrol):

You started with an equation, and now you only have an expression.

OpenStudy (anonymous):

idk, would it still = x/(x+8)

OpenStudy (anonymous):

alright, but i apparently did my math wrong somewhere cuz @agent0smith got (x+12)(x+8) = x(x+4)

OpenStudy (anonymous):

it looks more like u factored it

OpenStudy (agent0smith):

I didnt factor it... i just didn't expand it/multiply it out. Do it exactly as shown in this pic: http://www.mathsisfun.com/algebra/images/cross-multiply.gif

OpenStudy (agent0smith):

And show us what you get.

OpenStudy (agent0smith):

\[\Large \frac{( x+12) }{( x+4) } = \frac{ x }{ (x+8) }\]

OpenStudy (anonymous):

\[21x ^{2}+96 = x ^{2}+4x\]

OpenStudy (skullpatrol):

Work out the LHS: (x + 12)(x +8) using FOIL.

OpenStudy (anonymous):

yea i did thats what i got

OpenStudy (agent0smith):

How are you getting that left hand side? the right hand side is correct. I have no idea how you got the left, how are you getting 21x^2?.... immediately after cross multiplying, you should have \[\large (x+12)(x+8) = x( x+4)\]

OpenStudy (agent0smith):

Where does the 21 come from?

OpenStudy (anonymous):

when you multiply thos out you get x^2+8x+12x+96, then i added like terms

OpenStudy (agent0smith):

That's correct.. but where did the 21 come from...?

OpenStudy (agent0smith):

\[\large x^2+8x+12x+96 \] note that x^2 and x are *not* like terms. You can't add an x^2 term to an x term.

OpenStudy (anonymous):

8x + 12x = 20x then 20x + x^2 is 21x^2

OpenStudy (agent0smith):

20x + x^2 is 20x + x^2. Just like 2+x is 2+x, you can't combine unlike terms.

OpenStudy (anonymous):

alright so what your saying is that it should say: x^2 + 20x + 96 ?

OpenStudy (agent0smith):

correct

OpenStudy (anonymous):

ok i apparently forgot how do this, can u walk me through the next step?

OpenStudy (agent0smith):

Yes, you need to make sure you don't ever combine x^2 with x. You can't combine terms unless they're exactly the same, like you can't combine x^3 with x^2 for example.

OpenStudy (agent0smith):

\[\large x^2+20x+96 = x^2 +4x\]

OpenStudy (anonymous):

ok so now i put the constants one side?

OpenStudy (anonymous):

so then i would get 16x = 96?

OpenStudy (agent0smith):

Do it one step at a time. First, subtract x^2 from both sides \[\large x^2+20x+96 = x^2 +4x \] \[\large 20x+96 = 4x\] now subtract 4x from both sides.

OpenStudy (anonymous):

that would be -4x+20x+96=0

OpenStudy (agent0smith):

Correct. Now you can combine like terms, and then subtract 96 from both sides.

OpenStudy (anonymous):

it would be 16x=-96 then

OpenStudy (agent0smith):

Correct.

OpenStudy (anonymous):

then i divide to get -6?

OpenStudy (agent0smith):

Yep

OpenStudy (anonymous):

alright thanks you've been a huge help!

OpenStudy (agent0smith):

No prob :)